System Response

A LTIC system is specified by the equation

(D^2 + 5D + 6)y(t) = (D^2 + 7D + 11)x(t)

Find the zero-input response of the response y(t) and the impulse response h(t) if the initial conditions are y(0) = 0 and y'(0) = 1.
Answer #1

sol: Guiven LTIC system is given by (0+50+6) y(t) = (D+90+1)x(+ Initial conditions are yo) =0) { yo =) Zero Il Response (5+520, +26₂ = 0 -22,-36,-1 - C₂=1 [C2 =-11 1. Yo (t) = (ezt - p3t) / ® Impulse Response: we have hlt) = bn 8(t) at [P (D) yolt)Y you = kitk2=0 y(o) = -2k, - 3K2=1 Now wing initial Now sowing above two, we get K1=1 & K2=-1 Yn (t) = ezt _ eat P (1) Yo(there bo=bn . If impulse response of a system exist for t>0 then system is said to be causal.